# Analytic solutions for the Bianchi I universe coupled to several barotropic perfect fluids

### v2

@@ Abstract
We consider the Bianchi I geometry coupled to several species of comoving barotropic perfect fluids with a linear equation of state in the context of general relativity. The solution of the dynamics can be reduced to a quadrature, which can be explicitly performed in certain cases. In particular, we obtain the explicit solution for one species, as well as for two species, given their barotropic indices obey certain relation. These solutions include and generalize different models studied in the literature. For completeness, we analyze all the different possible signs of the matter energy densities, and we obtain a particularly interesting model in which an exotic species produces a bounce of the scale factor, providing a singularity-free cosmology, and then decays to leave a nonexotic component as the dominant fluid for large volumes.
@@ 1 Introduction
The paper is organized as follows. In Sec. 2 we present the basic variables and the Einstein equations for the Bianchi I geometry, with the matter content given by $n$ species of orthogonal (nontilted) barotropic perfect fluids. We also introduce the curvature invariants that we will later analyze for the exact solutions. Then, in Sec. 3, we fix the time gauge that will be used all along the paper, and give a general qualitative description of the dynamics based on an energy equation. Once this is set, in Sec. 4 we present analytic solutions for certain specific fluid components. More precisely, we obtain the solution for vacuum (Kasner), the general solution for one fluid species, as well as an analytic solution for two fluid species provided that their corresponding barotropic indices satisfy certain relation. In all the cases, we analyze the scenario with different signs of the energy densities, which, in particular, leads to a singularity-free cosmology for the model with two fluid species. Subsequently, in Sec. 5 the asymptotic behavior of the solutions is analyzed. Finally, in Sec. 6 we summarize and discuss the main results of the paper.
@@ 2.1 Basic variables and equations of motion
The Bianchi I geometry describes a spatially homogeneous, though anisotropic, universe. Its corresponding metric can be generically written as follows,
$ds^{2}=-N^{2}dT^{2}+a_{1}^{2}dx_{1}^{2}+a_{2}^{2}dx_{1}^{2}+a_{3}^{2}dx_{1}^{2},$ (2.1)
@@
which encode the spatial anisotropy of the universe. In particular, if all the scale factors are equal, $a_{1}=a_{2}=a_{3}$, then $\beta_{\pm}=0$. In addition, we also introduce the average scale factor
@@
where the prime $(^{\prime})$ denotes derivation with respect to the generic time $T$. Finally, the Einstein equations for this model can then be reduced to the following set of equations,
@@ 2.2 Equation of state
In summary, there are 3 geometric $(a,\beta_{+},\beta_{-})$ and $2n$ matter $(\rho_{s},p_{s})$ dynamical variables, with $s=1,\dots,n$, but, among the $(4+n)$ equations of motion — (2.7) and (2.8)–(2.11) — only $(3+n)$ are linearly independent. Therefore, as it is well known, in order to close the system, it is necessary to consider an equation of state for each fluid, which provides $n$ additional relations. In the following, we will assume barotropic fluids with $p_{s}=\omega_{s}\rho_{s}$. Different values of the constant $\omega_{s}$, known as the barotropic index, describe different matter contents of interest. For instance, $\omega_{s}=0$ corresponds to presureless dust, $\omega_{s}=1/3$ to relativistic particles, while $\omega_{s}=1$ defines stiff matter. This latter case is equivalent to a massless scalar field, for which the speed of sound equals the speed of light, and it corresponds to the maximum allowed value for $\omega_{s}$ that respects causality.
If one requires the dominant and strong energy conditions to be obeyed, $\rho_{s}$ must be nonnegative and $\omega_{s}$ is restricted to the interval $-\frac{1}{3}\leq\omega_{s}\leq 1$. However, there are some cases of interest that one can consider outside the commented ranges. In particular, a cosmological constant $\Lambda$ can be described as a fluid with $\omega_{s}=-1$, and with a contribution $\rho_{s}=\Lambda/\kappa$ to the energy density, which, if $\Lambda$ is negative, implies a negative $\rho_{s}$. Other more speculative fluids are also studied in the literature, like cosmic strings $(\omega_{s}=-1/3)$, domain-walls $(\omega_{s}=-2/3)$, or phantom energy $(\omega_{s}<-1)$, in some instances considering also their effects with negative energy density (see, e.g., Ref. [16] for a study in FLWR). Moreover, quantum-gravity effects are supposed to resolve the cosmological singularity, and, if one could describe those effects as an effective barotropic fluid, it would violate the energy conditions.
Therefore, in order to reproduce and generalize some results of the literature, we will not strictly impose the energy conditions, though we will exclude phantom energy with $\omega_{s}<-1$ and fluids that would violate causality with $\omega_{s}>1$ Thus, the sign of $\rho_{s}$ will be taken as arbitrary, and the $n$ species will be assumed to have different barotropic indices with $-1\leq\omega_{-}<\dots<\omega_{s}<\dots\omega_{+}\leq 1$ bounded by certain $\omega_{-}$ and $\omega_{+}$. These inequalities will be saturated in the case there is a cosmological constant in the model, and then $\omega_{-}=-1$, or if there is a stiff-matter species, and then $\omega_{+}=1$. For the case of a single component trivially $\omega_{-}=\omega_{+}$.
@@ 2.3 Curvature invariants
Therefore, the Weyl square is given as,
@@ 3 Gauge fixing and general description of the dynamics
Denoting with a dot the derivative with respect to the time $\tau$, in the $\tau$-gauge the equations of motion (2.8)–(2.11) read,
@@
where one has to take into account that $p=\sum_{s}\omega_{s}\rho_{s}$ and $\rho=\sum_{s}\rho_{s}$, with $\rho_{s}$ given in (2.12). The equations (3.3)–(3.4) do not depend on the matter content, and lead to a linear evolution of the shape parameters in the time $\tau$,
@@
with constants $k_{+},k_{-},p_{+},$ and $p_{-}$. The constants $k_{\pm}$ are pure gauge and can be absorbed in a redefinition of the coordinates, while $p_{\pm}$ are the canonical momenta of $\beta_{\pm}$, and completely encode the anisotropy of the model. In particular, the isotropic case corresponds to $p_{+}=0=p_{-}$.
@@
This is the equation of conservation of energy for a particle $a=a(\tau)$, with unit mass and zero energy, evolving under the effective potential $V(a)$. Analyzing the form of this potential, one can then infer the global evolution of $a$. In particular, since its total energy is zero, the particle can only move along regions where the potential $V(a)$ is nonpositive, while the roots of the potential, where $\dot{a}$ vanishes, represent boundaries of such regions. Before commenting the general qualitative behavior of the solutions in the $\tau$-gauge, let us first note that the vacuum and isotropic ($\rho=0=P)$ case is not dynamical, since the solution of the system is simply $a=a_{0}$. That is why, in some statements below, this degenerate $(\rho=0=P)$ case is excluded.
@@
On the one hand, if all densities $\rho_{0s}$ are positive or zero (excluding the case $\rho=0=P$), for all $a>0$ the potential $V(a)$ takes negative values, and it is a monotonically decreasing function of $a$. At $a=0$, $V(a)$ vanishes and, for $P\neq 0$, it has a maximum there, while, for the isotropic case $P=0$ (with $\rho\neq 0$), $a=0$ is an inflection point. This implies that, in either case, $a=0$ is not a turning point, rather it is an unstable equilibrium point that the system will only reach in an infinite amount of time. [^1] Note, however, that the amount of time is a gauge-dependent quantity. In a generic gauge with lapse $N$, the Friedmann equation (3.11) reads $\frac{{a^{\prime}}^{2}}{2}+\frac{N^{2}}{a^{6}}V(a)=0$. For instance, for the cosmological time $t$, $N=1$, and the corresponding effective potential is $V(a)/a^{6}$, which, instead of a maximum, presents a divergence at $a=0$. Therefore, in cosmic time, the singularity is reached in finite time as $\frac{da}{dt}$ tends to infinity. The image of the function $a=a(\tau)$ is the semi-infinite real line $(0,+\infty)$, and the velocity of the particle $\dot{a}$ increases with the value of $a$. Thus, choosing the outgoing (expanding) branch, the particle begins at $\tau\rightarrow-\infty$ at the origin $a=0$ with $\dot{a}=0$, and monotonically increases its velocity $\dot{a}$ as it moves to larger values of $a$.
@@
On the other hand, if some of the $\rho_{0s}$ are negative, the analysis is more involved since, in general, for $a>0$, the potential $V(a)$ will not be a monotonic function and it might have one (or several) roots. Such roots will bound allowed regions of $a$, and there might even exist more than one allowed region, defining different cosmological evolutions. In particular, below we present a detailed analysis for the case of two species with negative density, which, for certain ranges of the densities, describe two different cosmological evolutions. However, for this discussion, let us consider a simpler, though quite general, case: a model with negative cosmological constant $\Lambda<0$, for which $\omega_{\Lambda}=-1$ and $\rho_{0\Lambda}=\Lambda/\kappa<0$, while all the remaining species, if any, have $\rho_{0s}>0$ and $\omega_{s}\in(-1,1)$.
@@
In all the cases, the effect of the anisotropies is to lower the value of $V(a)$ with respect to its isotropic ($P=0$) counterpart. According to (3.11), this implies that, for any value of $a$, the velocity $|\dot{a}|$ will be larger than in the isotropic case. Moreover, in solutions with a recollapse at $a=a_{\rm max}$, the value of $a_{\rm max}$ will increase as $P$ increases, so that the recollapse will happen at larger volumes.
@@
From this expression, it is easy to check that it diverges towards the singularity $a\to 0$ in all the cases, except for the case with $P=0$, $\Lambda>0$, and no other matter content. In the latter case, $H\to{\rm sgn}(\dot{a})\sqrt{\Lambda/3}$ as $a\to 0$. Concerning large volumes, that is, as $a\to+\infty$, on the one hand, for monotonic solutions with $\Lambda\geq 0$, the Hubble factor tends to ${\rm sgn}(\dot{a})\sqrt{\Lambda/3}$. On the other hand, for recollapsing solutions, (with $\Lambda<0$, or some other exotic species with a negative density that triggers the recollapse) the image of $H(\tau)$ is the whole real line, and it vanishes at $a_{\rm max}$. In general, for a given value of $a$, the anisotropies increase the value of $|H|$ as compared to its isotropic counterpart. However, the effect of the anisotropies in $H$ dies off as $a$ expands, and under the presence of other species in the model with $\omega_{s}<1$, they will be subdominant for large volumes.
@@ 4 Explicit particular solutions
$a=a_{0}e^{P\tau},$ (4.2)
This is the well-known Kasner solution [1]. In order to see its usual form as a power-law in terms of the cosmological time, one can perform the change $t\to\tau$ by solving equation (3.1), that is,
@@
Moreover, by inverting the definition (2.2), and taking into account the evolution (3.6)–(3.7) of $\beta_{\pm}$, we can explicitly obtain the evolution of the different scale factors $a_{i}$ in terms of the cosmological time,
@@
Let us consider now the presence of only one species with a generic barotropic index $\omega\neq 1$ and $\rho_{0s}=\rho_{0}$. In this case, considering a positive density $\rho_{0}>0$, the solution reads as follows,
@@
with $\tau\in(-\infty,0)$. The isotropic case is included in the above expression as the limit $P\to 0$, and it explicitly reads
@@
This solution contains, as a particular case, the universe filled with dust $(\omega=0)$, which corresponds to the well-known Heckmann-Schücking [2] solution. This solution has been found in several (implicit) forms in the literature (see, for instance, Ref. [20]), but, to the best of our knowledge, it has not been given in this simple and explicit form, which we could obtain due to the gauge choice. Although the change to the cosmological time cannot be carried out explicitly for finite values of $\tau$, as will be detailed in the next section, where we will study the asymptotics of this solution, it is possible to do it for the limit of small and large volumes.
@@
with $\tau\in(-\infty,\infty)$. This solution contains an initial and a final singularity at $\tau\to-\infty$ and $\tau\to+\infty$, respectively, while $\tau=0$ corresponds to arecollapse. The domain of $a$ is thus bounded to $a\in(0,a_{\rm max}]$, where
@@
This case is a generalization of the models analyzed in Refs. [3, 4], where they considered a Bianchi I universe filled with stiff matter, dust and a cosmological constant $\Lambda$ (both positive and negative). As will be detailed below, these cases can be recovered from our results by simply imposing $\omega_{-}=-1$, $\omega_{+}=0$, $\rho_{0+}=\rho_{0{\rm dust}}$, $\rho_{0-}=\Lambda/\kappa$, and including the stiff-matter contribution by the replacement $P\to(P^{2}+\frac{\kappa}{3}\rho_{0{\rm stiff}})^{1/2}$. There are, however, other cases that may be of interest and are included in this analysis, for instance, a Bianchi I universe filled with radiation $(\omega_{1}=1/3)$ and cosmic strings $(\omega_{2}=-1/3)$.
For completeness, below we will present the explicit solution for all the possible signs of the densities. But, let us first comment several properties that can be inferred from the analysis of the potential $V(a)$. In particular, note that for small scales $a\to 0$, $\rho_{0+}$ will be dominant, while for large scales $a\to+\infty$, ‘$\rho_{0+}$ will be negligible and $\rho_{0-}$ will dominate. If both $\rho_{0+}$ and $\rho_{0-}$ are positive, the potential $V(a)$ does not have any positive root, it is negative all along $a\in[0,+\infty)$, and the model describes an indefinite cosmological expansion. But negative values of the densities $\rho_{0+}$ and $\rho_{0-}$ may introduce positive roots in $V(a)$, which, as commented above, are turning points that bound $a$. More precisely, $\rho_{0-}$ being the dominant species for $a\to+\infty$, if $\rho_{0-}<0$, this negative energy will always trigger a recollapse, which will bound $a$ from above, independently of the value of $\rho_{0+}$. That is, for $\rho_{0-}<0$ the potential $V(a)$ will have one, and only one, positive root $a_{\rm max}$. However, the case $\rho_{0+}<0$ and $\rho_{0-}>0$ is more involved and the existence of turning points depends on whether $\rho_{0-}$ exceeds certain threshold $\rho_{\rm threshold}:=\kappa\rho_{0+}^{2}/12P^{2}$. More precisely, if $\rho_{\rm threshold}<\rho_{0-}$, there are no positive roots of $V(a)$ and one gets an indefinite cosmological expansion. However, if the second species is not so energetic and $0\leq\rho_{0-}<\rho_{\rm threshold}$, $V(a)$ has two positive roots $a_{\pm}$, which leads to two different and independent cosmological evolutions. In one of them $a$ is bounded to $a\in(0,a_{-}]$ and it describes a recollapsing cosmology with an initial and a final singularity. In the other one, $a\in[a_{+},+\infty)$ is bounded from below, and it provides a singularity-free cosmology that begins with infinite volume, reaches a minimum at $a_{+}$, where it bounces back to expand then forever.
@@
On the other hand, for $\rho_{0+}<0$, any value of $\rho_{0-}$, and $P\neq 0$, there is a solution that can be written as
$a$ $=\left[{\frac{\kappa|\rho_{0+}|}{3P^{2}}\cosh^{2}\left({\frac{3}{2}(1-\omega_{+})P\tau}\right)}-\frac{\rho_{0-}}{|\rho_{0+}|}e^{3(1-\omega_{+})P\tau}\right]^{-\frac{1}{3(1-\omega_{+})}}.$ (4.14)
However, if $\rho_{0+}<0$ and $\rho_{0-}\in[0,\rho_{\rm threshold}]$, with $\rho_{\rm threshold}=\frac{\kappa\rho_{0+}^{2}}{12P^{2}}$, there is an additional solution of the form
$a$ $=\left[-{\frac{\kappa|\rho_{0+}|}{3P^{2}}\sinh^{2}\left({\frac{3}{2}(1-\omega_{+})P\tau}\right)}+\frac{\rho_{0-}}{|\rho_{0+}|}e^{3(1-\omega_{+})P\tau}\right]^{-\frac{1}{3(1-\omega_{+})}}.$ (4.15)
In the solutions (4.13) and (4.15) the isotropic case is included as the limit $P\to 0$, but (4.14) is not defined in such limit as this solution does not exist for $P=0$. Note that, if $P=0$ then $\rho_{\text{threshold}}\to+\infty$, and thus (4.15) is the only solution for $\rho_{0+}<0$, provided that $\rho_{0-}$ is positive.

### v3

@@ Abstract
We consider the Bianchi I geometry coupled to several species of comoving barotropic perfect fluids with a linear equation of state in the context of general relativity. The solution of the dynamics can be reduced to a quadrature, which can be explicitly performed in certain cases. In particular, we obtain the explicit solution for one species, as well as for two species, given their barotropic indices obey a certain relation. These solutions include and generalize different models studied in the literature. For completeness, we analyze all the different possible signs of the matter energy densities, and we obtain a particularly interesting model in which an exotic species produces a bounce of the scale factor, providing a singularity-free cosmology, and then decays to leave a nonexotic component as the dominant fluid for large volumes.
@@ 1 Introduction
The paper is organized as follows. In Sec. 2 we present the basic variables and the Einstein equations for the Bianchi I geometry, with the matter content given by $n$ species of orthogonal (nontilted) barotropic perfect fluids. We also introduce the curvature invariants that we will later analyze for the exact solutions. Then, in Sec. 3, we fix the time gauge that will be used all along the paper, and give a general qualitative description of the dynamics based on an energy equation. Once this is set, in Sec. 4 we present analytic solutions for certain specific fluid components. More precisely, we obtain the solution for vacuum (Kasner), the general solution for one fluid species, as well as an analytic solution for two fluid species provided that their corresponding barotropic indices satisfy a certain relation. In all of the cases, we analyze the scenario with different signs of the energy densities, which, in particular, leads to a singularity-free cosmology for the model with two fluid species. Subsequently, in Sec. 5 the asymptotic behavior of the solutions is analyzed. And, in Sec. 6 we summarize and discuss the main results of the paper. Finally, in the Appendix some plots for the model with two fluids are presented to illustrate the behavior of the different cases.
@@ 2.1 Basic variables and equations of motion
The Bianchi I geometry describes a spatially homogeneous, though anisotropic, universe. Its corresponding metric can be generically written as follows:
$ds^{2}=-N^{2}dT^{2}+a_{1}^{2}dx_{1}^{2}+a_{2}^{2}dx_{2}^{2}+a_{3}^{2}dx_{3}^{2},$ (2.1)
@@
which encode the spatial anisotropy of the universe. In particular, if all of the scale factors are equal, $a_{1}=a_{2}=a_{3}$, then $\beta_{\pm}=0$. In addition, we also introduce the average scale factor,
@@
where the prime $(^{\prime})$ denotes derivation with respect to the generic time $T$. Finally, the Einstein equations for this model can then be reduced to the following set of equations:
@@ 2.2 Equation of state
In summary, there are three geometric $(a,\beta_{+},\beta_{-})$ and $2n$ matter $(\rho_{s},p_{s})$ dynamical variables, with $s=1,\dots,n$, but, among the $(4+n)$ equations of motion (2.7)–(2.11) only $(3+n)$ are linearly independent. Therefore, as it is well known, in order to close the system, it is necessary to consider an equation of state for each fluid, which provides $n$ additional relations. In the following, we will assume barotropic fluids with $p_{s}=\omega_{s}\rho_{s}$. Different values of the constant $\omega_{s}$, known as the barotropic index, describe different matter contents of interest. For instance, $\omega_{s}=0$ corresponds to pressureless dust, $\omega_{s}=1/3$ to relativistic particles, while $\omega_{s}=1$ defines stiff matter. This latter case is equivalent to a massless scalar field, for which the speed of sound equals the speed of light, and it corresponds to the maximum allowed value for $\omega_{s}$ that respects causality.
If one requires the dominant and strong energy conditions to be obeyed, $\rho_{s}$ must be nonnegative and $\omega_{s}$ is restricted to the interval $-\frac{1}{3}\leq\omega_{s}\leq 1$. However, there are some cases of interest that one can consider outside the commented ranges. In particular, a cosmological constant $\Lambda$ can be described as a fluid with $\omega_{s}=-1$, and with a contribution $\rho_{s}=\Lambda/\kappa$ to the energy density, which, if $\Lambda$ is negative, implies a negative $\rho_{s}$. Other more speculative fluids are also studied in the literature, like cosmic strings $(\omega_{s}=-1/3)$, domain walls $(\omega_{s}=-2/3)$, or phantom energy $(\omega_{s}<-1)$, in some instances considering also their effects with negative energy density (see, e.g., Ref. [16] for a study in FLRW). Moreover, quantum-gravity effects are supposed to resolve the cosmological singularity, and, if one could describe those effects as an effective barotropic fluid, it would violate the energy conditions.
Therefore, in order to reproduce and generalize some results of the literature, we will not strictly impose the energy conditions, though we will exclude phantom energy with $\omega_{s}<-1$ and fluids that would violate causality with $\omega_{s}>1$ Thus, the sign of $\rho_{s}$ will be taken as arbitrary, and the $n$ species will be assumed to have different barotropic indices with $-1\leq\omega_{-}<\dots<\omega_{s}<\dots<\omega_{+}\leq 1$ bounded by certain $\omega_{-}$ and $\omega_{+}$. These inequalities will be saturated in the case there is a cosmological constant in the model, and then $\omega_{-}=-1$, or if there is a stiff-matter species, and then $\omega_{+}=1$. For the case of a single component trivially $\omega_{-}=\omega_{+}$.
@@ 2.3 Curvature invariants
Therefore, the Weyl square is given as
@@ 3 Gauge fixing and general description of the dynamics
Denoting with a dot the derivative with respect to the time $\tau$, in the $\tau$ gauge the equations of motion (2.8)–(2.11) read
@@
where one has to take into account that $p=\sum_{s}\omega_{s}\rho_{s}$ and $\rho=\sum_{s}\rho_{s}$, with $\rho_{s}$ given in (2.12). Equations (3.3) and (3.4) do not depend on the matter content, and lead to a linear evolution of the shape parameters in the time $\tau$,
@@
with constants $k_{+},k_{-},p_{+},$ and $p_{-}$. The constants $k_{\pm}$ are pure gauge and can be absorbed in a redefinition of the coordinates, [^1] Specifically, the change of coordinates to perform is $x_{1}\to x_{1}e^{-k_{+}-\sqrt{3}k_{-}}$, $x_{2}\to x_{2}e^{-k_{+}+\sqrt{3}k_{-}}$, and $x_{3}\to x_{3}e^{2k_{+}}$. while $p_{\pm}$ are the canonical momenta of $\beta_{\pm}$, and completely encode the anisotropy of the model. In particular, the isotropic case corresponds to $p_{+}=0=p_{-}$.
@@
This is the equation of conservation of energy for a particle $a=a(\tau)$, with unit mass and zero energy, evolving under the effective potential $V(a)$. Analyzing the form of this potential, one can then infer the global evolution of $a$. In particular, since its total energy is zero, the particle can only move along regions where the potential $V(a)$ is nonpositive, while the roots of the potential, where $\dot{a}$ vanishes, represent boundaries of such regions. Before commenting the general qualitative behavior of the solutions in the $\tau$ gauge, let us first note that the vacuum and isotropic ($\rho=0=P)$ case is not dynamical, since the solution of the system is simply $a=a_{0}$. That is why, in some statements below, this degenerate $(\rho=0=P)$ case is excluded.
@@
On the one hand, as shown in Fig. 2, if all densities $\rho_{0s}$ are positive or zero (excluding the case $\rho=0=P$), for all $a>0$ the potential $V(a)$ takes negative values, and it is a monotonically decreasing function of $a$. At $a=0$, $V(a)$ vanishes and, for $P\neq 0$, it has a maximum there, while, for the isotropic case $P=0$ (with $\rho\neq 0$), $a=0$ is an inflection point. This implies that, in either case, $a=0$ is not a turning point, rather it is an unstable equilibrium point that the system will only reach in an infinite amount of time. [^2] Note, however, that the amount of time is a gauge-dependent quantity. In a generic gauge with lapse $N$, the Friedmann equation (3.11) reads $\frac{{a^{\prime}}^{2}}{2}+\frac{N^{2}}{a^{6}}V(a)=0$. For instance, for the cosmological time $t$, $N=1$, and the corresponding effective potential is $V(a)/a^{6}$, which, instead of a maximum, presents a divergence at $a=0$. Therefore, in cosmic time, the singularity is reached in finite time as $\frac{da}{dt}$ tends to infinity. The image of the function $a=a(\tau)$ is the semi-infinite real line $(0,+\infty)$, and the velocity of the particle $\dot{a}$ increases with the value of $a$. Thus, choosing the outgoing (expanding) branch, the particle begins at $\tau\rightarrow-\infty$ at the origin $a=0$ with $\dot{a}=0$, and monotonically increases its velocity $\dot{a}$ as it moves to larger values of $a$.
Figure 1: Shape of $V(a)$, as defined in (3.10), for the specific scenario where all the densities $\rho_{0s}$ are positive or zero (excluding the case $\rho=0=P$), shown for both the isotropic $(P=0)$ and anisotropic $(P\neq 0)$ cases. As can be seen, in its domain $a>0$, $V(a)$ is a negative and monotonically decreasing function, with anisotropies ($P\neq 0$) making it more negative.
Figure 2: Shape of $V(a)$, as defined in (3.10), for the specific scenario of a negative cosmological constant with the remaining densities $\rho_{0s}$ positive, shown for both the isotropic and anisotropic cases. As can be seen, $V(a)$ has a root in $a_{\max}$, which increases with the anisotropies ($P\neq 0$). Moreover, the latter also decrease the value of $V(a)$.
@@
On the other hand, if some of the $\rho_{0s}$ are negative, the analysis is more involved since, in general, for $a>0$, the potential $V(a)$ will not be a monotonic function and it might have one (or several) roots. Such roots will bound allowed regions of $a$, and there might even exist more than one allowed region, defining different cosmological evolutions. In particular, below we present a detailed analysis for the case of two species with negative density, which, for certain ranges of the densities, describe two different cosmological evolutions.
@@
In all of the cases, the effect of the anisotropies is to lower the value of $V(a)$ with respect to its isotropic ($P=0$) counterpart. According to (3.11), this implies that, for any value of $a$, the velocity $|\dot{a}|$ will be larger than in the isotropic case. Moreover, in solutions with a recollapse at $a=a_{\rm max}$, the value of $a_{\rm max}$ will increase as $P$ increases, so that the recollapse will happen at larger volumes.
@@
From this expression, it is easy to check that it diverges towards the singularity $a\to 0$ in all the cases, except for the case with $P=0$, $\Lambda>0$, and no other matter content. In the latter case, $H\to{\rm sgn}(\dot{a})\sqrt{\Lambda/3}$ as $a\to 0$. Concerning large volumes, that is, as $a\to+\infty$, on the one hand, for monotonic solutions with $\Lambda\geq 0$, the Hubble factor tends to ${\rm sgn}(\dot{a})\sqrt{\Lambda/3}$. On the other hand, for recollapsing solutions (with $\Lambda<0$, or some other exotic species with a negative density that triggers the recollapse), the image of $H(\tau)$ is the whole real line, and it vanishes at $a_{\rm max}$. In general, for a given value of $a$, the anisotropies increase the value of $|H|$ as compared to its isotropic counterpart. However, the effect of the anisotropies in $H$ dies off as $a$ expands, and under the presence of other species in the model with $\omega_{s}<1$, they will be subdominant for large volumes.
@@ 4 Explicit particular solutions
$a=a_{0}e^{P\tau}.$ (4.2)
This is the well-known Kasner solution [1]. In order to see its usual form as a power law in terms of the cosmological time, one can perform the change $t\to\tau$ by solving Eq. (3.1), that is,
@@
Moreover, by inverting the definition (2.2), and taking into account the evolution (3.6) and (3.7) of $\beta_{\pm}$, we can explicitly obtain the evolution of the different scale factors $a_{i}$ in terms of the cosmological time,
@@
Let us consider now the presence of only one species with a generic barotropic index $\omega\neq 1$ and $\rho_{0s}=\rho_{0}$. In this case, considering a positive density $\rho_{0}>0$, the solution reads as follows,:
@@
with domain $\tau\in(-\infty,0)$. The form of this function is shown in Fig. 4 for different values of $\omega$. The isotropic case is included in the above expression as the limit $P\to 0$, and it explicitly reads
@@
This solution contains, as a particular case, the universe filled with dust $(\omega=0)$, which corresponds to the well-known Heckmann-Schücking [2] solution. Another particular case is the one presented in Ref. [20], involving an ekpyrotic fluid ($\omega=3$) with an additional stiff-matter component, which can be included in (4.9) by the replacement commented above $P\to(P+\frac{\kappa}{3}\rho_{0{\rm stiff}})^{1/2}$.
The solution (4.9) has been found in several (implicit) forms in the literature (see, for instance, Ref. [21]), but, to the best of our knowledge, it has not been given in this simple and explicit form, which we could obtain due to the gauge choice. Although the change to the cosmological time cannot be carried out explicitly for finite values of $\tau$, as will be detailed in the next section, where we will study the asymptotics of this solution, it is possible to do it for the limit of small and large volumes.
@@
with domain $\tau\in(-\infty,\infty)$. As can be seen in Fig. 4, this solution contains an initial and a final singularity at $\tau\to-\infty$ and $\tau\to+\infty$, respectively, while $\tau=0$ corresponds to a recollapse. The domain of $a$ is thus bounded to $a\in(0,a_{\rm max}]$, where
@@
Figure 3: Evolution of the average scale factor $a$ in terms of $\tau$ given by (4.9), for a single species and a positive density $\rho_{0}>0$. Different colors correspond to different barotropic indices $\omega$. As can be seen, for a given $\tau$, a larger value of $\omega$ implies a larger value of $a$.
Figure 4: Evolution of the average scale factor $a$ in terms of $\tau$ given by (4.11), for a single species and a negative density $\rho_{0}<0$. Different colors correspond to different barotropic indices $\omega$. As can be seen, for larger $\omega$, the $a$ is larger at each value of $\tau$. Hence, the corresponding maximum value $a_{\max}$, given by (4.12), is also larger.
@@
This case is a generalization of the models analyzed in Refs. [3, 4], where they considered a Bianchi I universe filled with stiff matter, dust and a cosmological constant $\Lambda$ (both positive and negative). As will be detailed below, these cases can be recovered from our results by simply imposing $\omega_{-}=-1$, $\omega_{+}=0$, $\rho_{0+}=\rho_{0{\rm dust}}$, $\rho_{0-}=\Lambda/\kappa$, and including the stiff-matter contribution by the replacement $P\to(P^{2}+\frac{\kappa}{3}\rho_{0{\rm stiff}})^{1/2}$. There are, however, other cases that may be of interest and are included in this analysis, for instance, a Bianchi I universe filled with radiation $(\omega_{+}=1/3)$ and cosmic strings $(\omega_{-}=-1/3)$.
For completeness, below we will present the explicit solution for all the possible signs of the densities. But, let us first comment several properties that can be inferred from the analysis of the potential $V(a)$. In particular, note that for small scales $a\to 0$, $\rho_{0+}$ will be dominant, while for large scales $a\to+\infty$, $\rho_{0+}$ will be negligible and $\rho_{0-}$ will dominate. If both $\rho_{0+}$ and $\rho_{0-}$ are positive, the potential $V(a)$ does not have any positive root, it is negative all along $a\in[0,+\infty)$, and the model describes an indefinite cosmological expansion (see Fig. 8). But negative values of the densities $\rho_{0+}$ and $\rho_{0-}$ may introduce positive roots in $V(a)$, which, as commented above, are turning points that bound $a$. More precisely, $\rho_{0-}$ being the dominant species for $a\to+\infty$, if $\rho_{0-}<0$, this negative energy will make $V(a)$ to take positive values at large values of $a$, independently of the value of $\rho_{0+}$. Thus, for $\rho_{0-}<0$, the potential $V(a)$ will have one, and only one, positive root $a_{\rm max}$. This value will bound $a$ from above, triggering a recollapse (see Fig. 8).
However, the case $\rho_{0+}<0$ and $\rho_{0-}>0$ is more involved and the existence of turning points depends on whether $\rho_{0-}$ exceeds certain threshold $\rho_{\rm threshold}:=\kappa\rho_{0+}^{2}/12P^{2}$. More precisely, if $\rho_{\rm threshold}<\rho_{0-}$, there are no real roots of $V(a)$ and one gets an indefinite cosmological expansion. However, if the second species is not so energetic and $0\leq\rho_{0-}<\rho_{\rm threshold}$, $V(a)$ has two positive roots $a_{\pm}$, which leads to two different and independent cosmological evolutions. In one of them $a$ is bounded to $a\in(0,a_{-}]$ and it describes a recollapsing cosmology with an initial and a final singularity. In the other one, $a\in[a_{+},+\infty)$ is bounded from below, and it provides a singularity-free cosmology that begins with infinite volume, reaches a minimum at $a_{+}$, where it bounces back to expand then forever. These two different scenarios are illustrated in Figs. 8 and 8, respectively.
@@
This evolution can be seen in Fig. 12. On the other hand, for $\rho_{0+}<0$, any value of $\rho_{0-}$, and $P\neq 0$, there is a solution that can be written as
$a$ $=\left[{\frac{\kappa|\rho_{0+}|}{3P^{2}}\cosh^{2}\left({\frac{3}{2}(1-\omega_{+})P\tau}\right)}-\frac{\rho_{0-}}{|\rho_{0+}|}e^{3(1-\omega_{+})P\tau}\right]^{-\frac{1}{3(1-\omega_{+})}},$ (4.14)
which is depicted in Fig. 12. However, if $\rho_{0+}<0$ and $\rho_{0-}\in[0,\rho_{\rm threshold}]$, with $\rho_{\rm threshold}=\frac{\kappa\rho_{0+}^{2}}{12P^{2}}$, there is an additional solution of the form
$a$ $=\left[-{\frac{\kappa|\rho_{0+}|}{3P^{2}}\sinh^{2}\left({\frac{3}{2}(1-\omega_{+})P\tau}\right)}+\frac{\rho_{0-}}{|\rho_{0+}|}e^{3(1-\omega_{+})P\tau}\right]^{-\frac{1}{3(1-\omega_{+})}},$ (4.15)
which is shown in Fig. 12. In the solutions (4.13) and (4.15) the isotropic case is included as the limit $P\to 0$, but (4.14) is not defined in such limit as this solution does not exist for $P=0$. Note that, if $P=0$ then $\rho_{\text{threshold}}\to+\infty$, and thus (4.15) is the only solution for $\rho_{0+}<0$, provided that $\rho_{0-}$ is positive.